A plane travesl 880 miles in 4 hours against the wind and 1000 miles in 2 hours with the same wind. Find the speed of the plane in still air and the speed of the wind.
Answer by mananth(16949)

let speed in still air be x
Let speed of wind be y
..
speed against wind = x-y
distance traveled = 880 miles
Time taken = 4 hours
speed = 880/4 = 220 mph
x-y =220
..
with wind 1000 miles
time = 2 hours
speed = 500 mph
x+y=500
..
Add the two equations
x-y+x+y=720
2x=720
x=360 mph speed of plane in still air
x-y=220
plug the value of x in the equation
y=140 mph wind speed